**Please include solver

2. A steel mill produces two types of steel alloy: boral and chromal. Production of each alloy requires three processes: Box anneal, Cold Roll, and Strand anneal.

Production capacities are:

Box anneal: 4,000 hours/month

Cold Roll: 500 hours/month

Strand anneal: 1,000 hours/month

Production rates in tons:

Box anneal

Cold roll pass 1

Stand anneal

Cold roll pass 2

Boral

4

72

11

36

Chromal

2

Not required

20

24

The maximum demand for boral is 10,500 tons/month and for chromal 6,000 tons/month. The contributions/ton are boral: $25, and chromal: $35. What combination of boral and chromal maximizes total monthly contribution?
Answer #1

Number of Boral=X1

Number of Chromal=X2

Objective Functions= 25*X1+35*X2

Constraints:

X1=<10500

X2=<6000

4*X1+2*X2=<4000

108*X1+24*X2=<500

11*X1+20*X2=<1000

X1>=0

X2>=0

Hence, combination of boral and chromal maximizes total monthly contribution= (0,20.83)

1 B C D E F G
2 contributions/ton for Boral 25
3 contributions/ton for Chromal 35
4
5 Boral Chromal
6 Ton 0.00 20.83
7 contributions/ton 25 35
8 Boral Chromal
9 Box Anneal 4 2 41.66667 =< 4000
10 Cold Roll 108 24 500 =< 500
11 Stand Anneal 11 20 416.6667 =< 1000
12 1 0 0 =< 10500
13 0 1 20.83333 =< 6000
14 1 0 0 >= 0
15 0 1 20.83333 >= 0
16 Total Contributions 729.17

Solver Parameter:

Excel Formula:

1 B C D E F G
2 contributions/ton for Boral 25
3 contributions/ton for Chromal 35
4
5 Boral Chromal
6 Ton 0 20.8333333333333
7 contributions/ton 25 35
8 Boral Chromal
9 Box Anneal 4 2 =SUMPRODUCT($C$6:$D$6,C9:D9) =< 4000
10 Cold Roll 108 24 =SUMPRODUCT($C$6:$D$6,C10:D10) =< 500
11 Stand Anneal 11 20 =SUMPRODUCT($C$6:$D$6,C11:D11) =< 1000
12 1 0 =SUMPRODUCT($C$6:$D$6,C12:D12) =< 10500
13 0 1 =SUMPRODUCT($C$6:$D$6,C13:D13) =< 6000
14 1 0 =SUMPRODUCT($C$6:$D$6,C14:D14) >= 0
15 0 1 =SUMPRODUCT($C$6:$D$6,C15:D15) >= 0
16 Total Contributions =SUMPRODUCT(C6:D6,C7:D7)